• A person can do a job in 5hrs, working with another person the job take 3hrs, How long will it take the second person to do the job?
Answer by nerdybill(7384)

A person can do a job in 5hrs, working with another person the job take 3hrs, How long will it take the second person to do the job?
.
Let x = time (hours) it takes the second person
then
3(1/x + 1/5) = 1
multiplying both sides by 5x:
3(5 + x) = 5x
15 + 3x = 5x
15 = 2x
15/2 = x
7.5 hours = x